# DP-CDA: An Algorithm for Enhanced Privacy Preservation in Dataset Synthesis Through Randomized Mixing

### v1

@@ 1 Introduction
In the realm of machine learning or deep learning applications, it is widely acknowledged that the dataset is of utmost importance, as it contains sensitive information like medical records, personal photos/information, or proprietary data during the training process. Organizations, including private companies, government agencies, and hospitals, routinely handle substantial volumes of personal information about their clients, customers, or patients. This data is often private or sensitive in nature, underscoring the importance of implementing systems and methodologies that can effectively analyze the data while upholding privacy and ensuring the confidentiality of individuals. It should be noted that using standard machine learning techniques to train a model can potentially lead to data breaches [1]. Therefore, a prudent approach involves considering the creation of a “distilled version” of the original dataset, and subsequently training the model exclusively using that version to ensure the protection of sensitive information [2, 3, 4]. Note that, simple anonymization does not provide any privacy in the presence of auxiliary information [5]. More specifically, it’s important to be aware of “composition attacks”, where malicious actors can combine algorithm output with other information to identify individuals in the dataset [6].
@@
Related Works. Several differentially private data publishing algorithms have been proposed over the years [9, 10]. Zhu et al. conducted a comprehensive analysis of these algorithms, offering a clear and detailed overview [3]. However, these algorithms face challenges when scaling with larger data sets. That is, the computational complexity increases significantly with high-dimensional data, often exhibiting exponential growth [8]. As a result, many of these algorithms can become impractical for deep learning applications, particularly those that involve high-dimensional datasets.
@@
There is a considerable amount of work on the application of differentially private algorithms to various machine learning problems, including regression [17, 18, 19], online learning [20], graphical models [21], empirical risk minimization [22, 23, 24, 22, 5, 25], and deep learning [26, 27, 28]. The objectives of these algorithms differ slightly from those of the data publishing algorithms mentioned earlier. While differentially private data publishing algorithms focus on designing differentially private datasets at the input end of the pipeline, differentially private algorithms aim to create learning algorithms that yield differentially private models at the output end. Since any model trained on a differentially private dataset is also differentially private [7], the goals of these learning algorithms represent a more relaxed version of the data publishing algorithms. Unfortunately, these differentially private algorithms do not always guarantee data privacy in adverse scenarios. If an attacker gains access to the input side of the pipeline, the entire algorithm may be compromised, thus undermining the security of user data.
@@
In the past few years, several methods have been proposed for differentially private data publishing algorithms using mixtures. Karakus et al. demonstrated that a simple linear model can be trained with noiseless mixtures [29]. The exploration of learning nonlinear models with mixtures has also been addressed in previous studies [30, 31, 32, 33, 34]. For instance, Tokozune et al. demonstrated that sound recognition models can be trained using mixtures of audio signals [30]. Similarly, other studies have indicated that it is feasible to train an image classification model using image mixtures [31, 32, 33]. However, these studies primarily focused on noiseless mixtures involving only two or three data points. Note that, Lee et al. eventually proved that it is possible to train deep neural networks with a high order of mixture while incorporating added noise to enhance privacy [8].
@@
Recently, data publishing algorithms have been automated by incorporating a deep learning framework  [35]. The integration of a generative adversarial network for producing differentially private datasets has gained much popularity  [36],  [37]. Although GAN-generated synthetic images offer differential privacy (DP) guarantees and maintain the utility of the original private images, they often appear visually similar to those private images. This similarity poses a challenge in fulfilling a key privacy requirement, which is the need for visual dissimilarity. Furthermore, training a GAN can be difficult due to several issues, including vanishing gradients, mode collapse, training instability, and convergence failures. In this work, we consider the problem of releasing synthetic data for training an image classification model, while ensuring $(\epsilon,\delta)$-differential Privacy (DP) guarantees and providing utility close to that of conventional DP training. We note that the synthetic data generating mechanism is computationally simple, and can be adapted for high-dimensional data.
@@ 2 Background and Problem Formulation
We review some definitions, theorems, and propositions that are necessary for our problem formulation according to [5, 38, 25].
@@
Definition 1 (($\epsilon,\delta$)-DP [39]).
[…]
Definition 2 ($\ell_{2}$ sensitivity [39]).
Definition 3 (Gaussian Mechanism [40]).
Definition 4 (Rényi Differential Privacy (RDP) [42]).
Proposition 1 (From RDP to DP [42]).
Proposition 2 (Composition of RDP [42]).
Proposition 3 (RDP and Gaussian Mechanism [42]).
@@ 4 Experimental Results
Note that, our model includes batch normalization and dropout to enhance training stability and reduce overfitting, i.e., robust for image classification tasks. We utilized the Adam optimizer with an initial learning rate of 0.001, and employed a learning rate scheduler to adjust the learning rate as needed. As mentioned before, we recorded the accuracies on the original test partition of the base dataset, which is unmixed and noise-free.
@@ 5 Conclusion
In this work, we propose DP-CDA, a synthetic data generation algorithm based on the works of [8], and perform a tighter accounting of the privacy guarantee. DP-CDA combines randomly selected data points from a specific class and induces additive noise for preserving privacy. We empirically demonstrate that given a particular dataset, the number of data-points utilized for generating each synthetic data point has an optimal value, which offers the best utility and dictates the overall privacy budget. Through our extensive experiments, we show that DP-CDA significantly outperforms the existing data publishing algorithms, while being less computationally complex, providing a strong privacy guarantee. An interesting direction for future work would be to incorporate DP-CDA in federated settings, and to perform a theoretical analysis for choosing the optimal order of mixture.
@@ Acknowledgement
The authors extend their thanks to the authorities of the Department of Electrical and Electronic Engineering and Bangladesh University of Engineering and Technology (BUET) for providing constant support throughout this research.

### v2

@@ 1 Introduction
In the realm of machine learning or deep learning applications, it is widely acknowledged that the dataset is of utmost importance, as it contains sensitive information like medical records, personal photos/information, or proprietary data during the training process. Organizations, including private companies, government agencies, and hospitals, routinely handle substantial volumes of personal information about their clients, customers, or patients. This data is often private or sensitive in nature, underscoring the importance of implementing systems and methodologies that can effectively analyze the data while upholding privacy and ensuring the confidentiality of individuals. It should be noted that using standard machine learning techniques to train a model can potentially lead to data breaches [1]. Therefore, a prudent approach involves considering the creation of a “distilled version” of the original dataset, and subsequently training the model exclusively using that version to ensure the protection of sensitive information [2, 3, 4]. It should be noted here that only anonymization can not provide any privacy in the presence of auxiliary information [5]. More specifically, it’s important to be aware of “composition attacks”, where malicious actors can combine algorithm output with other information to identify individuals in the dataset [6].
@@
Related Works. Over the years, numerous algorithms have been thoughtfully developed for differentially private data publishing [9, 10]. Zhu et al. conducted a comprehensive analysis of these algorithms, offering a clear and detailed overview [3]. However, these algorithms face challenges when scaling with larger data sets. That is, the computational complexity increases significantly with high-dimensional data, often exhibiting exponential growth [8]. As a result, many of these algorithms can become impractical for deep learning applications, particularly those that involve high-dimensional datasets.
@@
There is a considerable amount of work on the application of differentially private algorithms to various machine learning problems, including regression [17, 18, 19], online learning [20], graphical models [21], empirical risk minimization [22, 23, 24, 22, 5, 25], and deep learning [26, 27, 28]. The objectives of these algorithms differ slightly from those of the data publishing algorithms mentioned earlier. While differentially private data publishing algorithms focus on designing differentially private datasets at the input end of the pipeline, differentially private algorithms aim to create learning algorithms that yield differentially private models at the output end. Any model trained on a dataset that is differentially private will also maintain differential privacy. [7], the goals of these learning algorithms represent a more relaxed version of the data publishing algorithms. Unfortunately, these differentially private algorithms do not always guarantee data privacy in adverse scenarios. If an attacker gains access to the input side of the pipeline, the entire algorithm may be compromised, thus undermining the security of user data.
@@
In the past few years, several methods have been proposed for differentially private data publishing algorithms using mixtures. Karakus et al. illustrated that it is possible to effectively train a straightforward linear model by utilizing mixtures that are free from noise [29]. The exploration of learning nonlinear models with mixtures has also been addressed in previous studies [30, 31, 32, 33, 34]. For instance, Tokozune et al. demonstrated that sound recognition models can be trained using mixtures of audio signals [30]. Similarly, other studies have indicated that it is feasible to train an image classification model using image mixtures [31, 32, 33]. However, these studies primarily focused on noiseless mixtures involving only two or three data points. Note that, Lee et al. eventually proved that it is possible to train deep neural networks with a high order of mixture while incorporating added noise to enhance privacy [8].
@@
Recently, data publishing algorithms have been automated by incorporating a deep learning framework  [35]. The integration of a generative adversarial network for producing differentially private datasets has gained much popularity  [36],  [37]. Although GAN-generated synthetic images offer differential privacy (DP) guarantees and maintain the utility of the original private images, they often appear visually similar to those private images. This similarity poses a challenge in fulfilling a key privacy requirement, which is the need for visual dissimilarity. Furthermore, training a GAN can be difficult due to several issues, including vanishing gradients, mode collapse, training instability, and convergence failures. In this work, we consider the problem of releasing synthetic data for training an image classification model, while ensuring $(\epsilon,\delta)$-Differential Privacy (DP) guarantees and providing utility close to that of conventional DP training. We note that the synthetic data-generating mechanism is computationally simple, and can be adapted for high-dimensional data.
@@ 2 Background and Problem Formulation
We reviewed some definitions, theorems, and propositions, which are necessary for our problem formulation, according to [5, 38, 25] in the supplementary materials (A).
@@ 4 Experimental Results
Note that, our model includes batch normalization and dropout to enhance training stability and reduce overfitting, i.e., robust for image classification tasks. We chose to implement the Adam optimizer, setting an initial learning rate of 0.001 and employed a learning rate scheduler to adjust the learning rate as needed. As mentioned before, we recorded the accuracies on the original test partition of the base dataset, which is unmixed and noise-free.
@@ 5 Conclusion
In this work, we propose DP-CDA, a synthetic data generation algorithm based on the works of [8], and perform a tighter accounting of the privacy guarantee. DP-CDA combines randomly selected data points from a specific class and induces additive noise for preserving privacy. We empirically demonstrate that given a particular dataset, the number of data points utilized for generating each synthetic data point has an optimal value, which offers the best utility and dictates the overall privacy budget. Through our extensive experiments, we show that DP-CDA significantly outperforms the existing data publishing algorithms, while being less computationally complex, providing a strong privacy guarantee. An interesting direction for future work would be to incorporate DP-CDA in federated settings and to perform a theoretical analysis for choosing the optimal order of mixture.
@@ Acknowledgement
The authors express their gratitude to the authorities of the Department of Electrical and Electronic Engineering at Bangladesh University of Engineering and Technology (BUET) for their continuous support during this research.
Appendix A Relevant Definitions and Theorems
Definition 1 (($\epsilon,\delta$)-DP [44]).
[…]
Definition 2 ($\ell_{2}$ sensitivity [44]).
Definition 3 (Gaussian Mechanism [45]).
Definition 4 (Rényi Differential Privacy (RDP) [47]).
Proposition 1 (From RDP to DP [47]).
Proposition 2 (Composition of RDP [47]).
Proposition 3 (RDP and Gaussian Mechanism [47]).